OpenStudy (anonymous):

find the intercepts of the polynomial x^4-9x^2

OpenStudy (amoodarya):

x^4-9x^2 =x^2(x^2-9) =x*x*(x-3)(x+3) can you go on ?

OpenStudy (anonymous):

yes! I had the x^2(x^2-9) I just wasn't sure if that was right. So I factor the x^2-9 to get the (x-3)(x+3) correct?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so how do you get the point of -20 when graphing?

OpenStudy (anonymous):

the y is -20 when x is -2

OpenStudy (anonymous):